A patient is prescribed sotalol (Betapace) 320 mg orally once a day. Sotalol comes in 80 mg tablets. How many tablets would you give for each daily dose?______tablets

You need to administer digoxin (Lanoxin) 0.25 mg PO QD. You have digoxin (Lanoxin) 0.125 mg tablets, How many tablets should you administer? ______tablets

A patient who weighs 220 lb is to receive labetalol 0.25 mg/kg intravenously. The solution of labetalol on hand is 100 mg/20 mL. How many milliliters of labetalol will you prepare for a correct dose?______mL
Answer #1

ANSWER.

* Calculation of number of tablets required.

         Ordered dose = 320 mg

         Available dose = 80 mg/tablet.

   Number of tablets required = Ordered dose / Available dose

                                               = 320/80

                                               = 4 tablets.

* Calculation of number of tablets required.

          Ordered dose = 0.25 mg.

          Available dose = 0.125 mg/tablet.

Number of tablets required = Ordered dose /Available dose

                                               = 0.25/0.125

                                               = 2 tablets PO QD

* Calculation of volume(ml) of drug required.

          Formula

             Volume(ml) = Desired dose /Dose in hand * Quantity.

      Here

        Ordered dose = 0.25 mg/kg

        Weight of the patient = 220 lb

     Convert lb to kg for calculation

                 1 lb = 0.454 kg

        Weight of the patient = 220 * 0.454

                                          = 99.8 kg.

         So

                 Desired dose (D) = 0.25 * 99.8

                                              = 24.95 mg.

                 Dose in hand (H) = 100 mg

                    Quantity (Q) = 20 ml

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 24.95 / 100 * 20

                                                        = 0.25 * 20

                                                        = 5 ml.